OpenStudy (anonymous):

If f is continuously differentiable function and lim f(x)+f'(x)=L x--> inf prove lim f(x)=L x-->inf

OpenStudy (perl):

I can think of an example where this is true. Let \(\bf f(x) = e^{-x}\) , then \(\bf f'(x) = -e^{-x}\)

OpenStudy (perl):

$$ \Large{ \lim_ {x \to \infty} e^{-x} + (-e^{-x}) = 0 - 0 = 0 } $$

OpenStudy (perl):

we can make it more interesting

OpenStudy (perl):

$$ \Large \rm{ Suppose \\ f(x) = e^{-x} + k \\ f ' (x) = -e^{-x} \\ } \Large{ Then \\ \lim_ {x \to \infty} ~ f(x) + f ' (x) \\ =\lim_ {x \to \infty} (e^{-x}+k) + (-e^{-x}) \\ = (e^{-\infty }+k)+ (- e^{-\infty } ) \\= (\frac{1}{e^{\infty }}+ k) + (-\frac{1}{e^{\infty }}) \\~\\ = (0 + k ) + (-0) = k } $$

OpenStudy (perl):

thats not a proof, that was a plausibility check for the statement. it looks like its true.

OpenStudy (perl):

your statement is equivalent to proving that lim(x->oo) f ' (x) = 0

OpenStudy (perl):

the only way for lim f ' (x) to converge as x-> infinity, is for the limit to go to zero

OpenStudy (perl):



OpenStudy (perl):

otherwise the limit will oscillate or diverge

OpenStudy (perl):